# arXiv:2603.02252v1[eess.AS] 27 Feb 2026

### v1

@@ Abstract
We introduce Whisper-RIR-Mega, a benchmark dataset of paired clean and reverberant speech for evaluating automatic speech recognition (ASR) robustness to room acoustics. Each sample pairs a clean LibriSpeech utterance with the same utterance convolved with a real room impulse response from the RIR-Mega corpus, with stratified splits by reverberation time (RT60) and direct-to-reverberant ratio (DRR). We evaluate five Whisper models (tiny through large-v3) on 1600 test samples and report word error rate (WER) and character error rate (CER) under clean and reverberant conditions. Reverberation consistently degrades performance across all model sizes; the reverb penalty in WER ranges from 0.12 to 1.07 percentage points depending on the model. We release the dataset, evaluation code, and baseline results to support reproducible research on robust ASR.
@@ 1 Introduction
We run baseline evaluations using five OpenAI Whisper models (tiny, base, small, medium, large-v3) on 1600 test samples. All models show higher WER and CER under reverberant speech; the smallest model (tiny) exhibits the largest reverb penalty in WER, while small and medium show the smallest penalty. We release the dataset on Hugging Face, along with evaluation code and a leaderboard, to encourage reproducible benchmarks and further work on robust ASR.
@@ 2 Related Work
Reverberation robust ASR has been addressed via multi-condition training, dereverberation front-ends, and end-to-end systems. Benchmark datasets include REVERB [2], CHiME, and others that provide simulated or real reverberant speech. RIR-Mega [4] is a large-scale simulated RIR dataset with machine-friendly metadata (RT60, DRR, C50); RIR-Mega-Speech [5] builds a reverberant speech corpus from LibriSpeech and RIR-Mega with per-file acoustic annotations and reproducible evaluation. Whisper [6] provides strong open ASR baselines. Our benchmark combines LibriSpeech and RIR-Mega [4] in a paired clean-reverb design with stratified splits and public baselines, complementary to [5].
@@ 3.1 Speech and RIR Sources
We use LibriSpeech test-clean (16 kHz) as the speech source. Each utterance is paired with a single room impulse response from the RIR-Mega dataset. RIR-Mega provides measured RIRs with metadata including RT60 (reverberation time) and DRR (direct-to-reverberant ratio). We use the train split of RIR-Mega and sample one RIR per utterance. When RT60 (or DRR) metadata is available, we stratify sampling across quantile bins so that the resulting dataset is balanced across acoustic conditions.
@@ 4.1 Models
We evaluate five Whisper models: openai/whisper-tiny, base, small, medium, and
openai/whisper-large-v3. Decoding uses beam size 5, best-of 5, temperature 0, and language set to English. All runs are performed on CPU with FP16 disabled for reproducibility.
@@ 4.2 Metrics
We report word error rate (WER) and character error rate (CER) using the jiwer library. For each model we compute mean WER and mean CER over the test set under two conditions: clean (original utterance) and reverb (convolved utterance). The reverb penalty is defined as the difference (reverb minus clean) in WER or CER.
@@ 5 Results
Table 1 reports mean WER and CER for each model on clean and reverberant test sets (1600 samples each condition). All models achieve similar clean WER (about 98.6–101.4%); the high values are consistent with the fact that we do not apply any normalization to the reference (e.g. casing or punctuation), so differences in capitalization and punctuation between reference and hypothesis can inflate WER. The relative comparison across conditions and models remains valid.
Reverberation increases WER for every model. The reverb penalty in WER (delta WER) is largest for Whisper-tiny (1.07 percentage points) and smallest for Whisper-small (0.12) and Whisper-medium (0.15). Whisper-base and Whisper-large-v3 show intermediate penalties (0.58 and 0.54 percentage points). CER follows a similar pattern: reverb degrades performance, with the smallest penalty for small and medium and the largest for tiny.
Figure 1 illustrates clean versus reverberant WER per model. When RIR metadata (RT60, DRR) is available in the dataset, we also produce plots of WER versus RT60 bin and WER versus DRR bin (reverb condition only); these figures are included in the repository and on the benchmark Space.
@@
Whisper-tiny 99.30 100.36 1.07 83.47 84.31 0.84
Whisper-base 98.83 99.41 0.58 83.33 83.70 0.37
Whisper-small 98.60 98.72 0.12 83.24 83.27 0.04
Whisper-medium 98.60 98.74 0.15 83.24 83.21 -0.02
Whisper-large-v3 101.36 101.90 0.54 84.94 85.91 0.97
@@ 6 Discussion
The results show that reverberation hurts ASR performance across all Whisper sizes, but the amount of degradation varies. Smaller models (tiny) are more sensitive to reverb in terms of WER delta; medium and small show the smallest penalty. This suggests that model capacity and training objectives interact with acoustic degradation in a non-trivial way and that robustness should be evaluated explicitly on benchmarks like Whisper-RIR-Mega.
@@ 7 Conclusion
We introduced Whisper-RIR-Mega, a paired clean-reverberant speech benchmark built from LibriSpeech and RIR-Mega, with stratified splits and baseline results for five Whisper models. Reverberation consistently increases WER and CER, with the reverb penalty varying by model size. The dataset, code, and leaderboard are publicly available to support reproducible research on robust ASR.
@@
The dataset is hosted at https://huggingface.co/datasets/mandipgoswami/whisper-rirmega-bench. Evaluation code and reproduction instructions are in the repository https://github.com/mandip42/whisper-rirmega-bench. An interactive Space with the leaderboard and submission interface is at https://huggingface.co/spaces/mandipgoswami/whisper-rirmega-benchmark.

### v3

# Whisper-RIR-Mega: A Paired Clean-Reverberant Speech Benchmark for ASR Robustness to Room Acoustics

@@ Abstract
We introduce Whisper-RIR-Mega, a benchmark dataset of paired clean and reverberant speech for evaluating automatic speech recognition (ASR) robustness to room acoustics. Each sample pairs a clean LibriSpeech utterance with the same utterance convolved with a real room impulse response from the RIR-Mega corpus, with stratified splits by reverberation time (RT60) and direct-to-reverberant ratio (DRR). We evaluate five Whisper models (tiny through large-v3) on 1600 test samples and report word error rate (WER) and character error rate (CER) under clean and reverberant conditions. Reverberation consistently degrades performance across all model sizes; the reverb penalty in WER ranges from 2.31 to 15.50 percentage points depending on the model. Whisper-large-v3 shows the smallest penalty; Whisper-tiny shows the largest. We release the dataset, evaluation code, and baseline results to support reproducible research on robust ASR.
@@ 1 Introduction
We run baseline evaluations using five OpenAI Whisper models (tiny, base, small, medium, large-v3) on 1600 test samples. All models show higher WER and CER under reverberant speech; the smallest model (tiny) exhibits the largest reverb penalty in WER (15.50 percentage points), while large-v3 shows the smallest (2.31 percentage points). We release the dataset on Hugging Face, along with evaluation code and a leaderboard, to encourage reproducible benchmarks and further work on robust ASR.
@@ 2 Related Work
Reverberation robust ASR has been addressed via multi-condition training, dereverberation front-ends, and end-to-end systems. Benchmark datasets include the REVERB challenge [2], CHiME [1], and others that provide simulated or real reverberant speech. LibriSpeech [3] is widely used as a clean-speech source. RIR-Mega [4] is a large-scale simulated RIR dataset with machine-friendly metadata (RT60, DRR, C50); RIR-Mega-Speech [5] builds a reverberant speech corpus from LibriSpeech and RIR-Mega with per-file acoustic annotations and reproducible evaluation. Whisper [6] provides strong open ASR baselines. Our benchmark combines LibriSpeech [3] and RIR-Mega [4] in a paired clean-reverb design with stratified splits and public baselines, complementary to [5].
@@ 3.1 Speech and RIR Sources
We use LibriSpeech test-clean [3] (16 kHz) as the speech source. Each utterance is paired with a single room impulse response from the RIR-Mega dataset. RIR-Mega provides measured RIRs with metadata including RT60 (reverberation time) and DRR (direct-to-reverberant ratio). We use the train split of RIR-Mega and sample one RIR per utterance. When RT60 (or DRR) metadata is available, we stratify sampling across quantile bins so that the resulting dataset is balanced across acoustic conditions.
@@ 4.1 Models
We evaluate five Whisper models: openai/whisper-tiny, base, small, medium, and openai/whisper-large-v3. Decoding uses beam size 5, best-of 5, temperature 0, and language set to English. All runs are performed on CPU with FP16 disabled for reproducibility.
@@ 4.2 Metrics
We report word error rate (WER) and character error rate (CER) using the jiwer library with standard normalization (lowercase, punctuation removal, whitespace collapse). For each model we compute mean WER and mean CER over the test set under two conditions: clean (original utterance) and reverb (convolved utterance). The reverb penalty is defined as the difference (reverb minus clean) in WER or CER.
@@ 5 Results
Table 1 reports mean WER and CER for each model on clean and reverberant test sets (1600 samples each condition). WER and CER use standard normalization (lowercase, punctuation removal). Clean WER ranges from 29.0% (Whisper-large-v3) to 54.9% (Whisper-tiny).
Reverberation increases WER for every model. The reverb penalty in WER (delta WER) is largest for Whisper-tiny (15.50 percentage points) and smallest for Whisper-large-v3 (2.31 percentage points). Whisper-medium and Whisper-small show intermediate penalties (5.94 and 7.44 percentage points). CER follows a similar pattern: reverb degrades performance, with the smallest penalty for Whisper-medium (0.48 pp) and the largest for Whisper-tiny (3.80 pp).
Figure 1 illustrates clean versus reverberant WER per model. When RIR metadata (RT60, DRR) is available in the dataset, the pipeline also produces WER versus RT60 bin and WER versus DRR bin (reverb condition); those plots are included in the repository and on the benchmark Space.
@@
Whisper-tiny 54.88 70.38 15.50 3.72 7.52 3.80
Whisper-base 46.50 57.94 11.44 3.08 4.88 1.80
Whisper-small 35.88 43.31 7.44 2.05 3.04 0.99
Whisper-medium 30.06 36.00 5.94 1.90 2.38 0.48
Whisper-large-v3 29.00 31.31 2.31 3.07 4.51 1.44
@@ 6 Discussion
The results show that reverberation hurts ASR performance across all Whisper sizes, but the amount of degradation varies. Smaller models (tiny) are more sensitive to reverb in terms of WER delta; Whisper-large-v3 shows the smallest penalty. The relationship between model size and reverb sensitivity is monotonic: larger models exhibit smaller WER degradation. Robustness should be evaluated explicitly on benchmarks like Whisper-RIR-Mega.
@@ 7 Conclusion
We introduced Whisper-RIR-Mega, a paired clean-reverberant speech benchmark built from LibriSpeech and RIR-Mega, with stratified splits and baseline results for five Whisper models. Reverberation consistently increases WER and CER; the reverb penalty ranges from 2.31 to 15.50 percentage points in WER, with larger models showing greater robustness. The dataset, code, and leaderboard are publicly available to support reproducible research on robust ASR.
@@
The dataset is hosted at https://huggingface.co/datasets/mandipgoswami/whisper-rirmega-bench. Evaluation code and reproduction instructions are in the repository https://github.com/mandipgoswami/Whisper_RIRMega. An interactive Space with the leaderboard and submission interface is at https://huggingface.co/spaces/mandipgoswami/whisper-rirmega-benchmark.